In a flexible-walled container, 300 L of gas is prepared at 600 mm Hg and 200 oC. The gas is placed in a tank under high pressure. When the tank cools to 20 oC, the pressure of the gas is 30 atm. What is the new volume of the gas?



1
Expert's answer
2021-02-22T05:54:23-0500

Initial Pressure "P_1" = 600mm Hg = 600/760 atm = 0.7894 atm

Initial volume "V_1" = 300L

Initial temperature "T_{1}" = 473K

Final temperature "T_2" = 293K

Final Pressure "P_2" = 30 atm

Since, moles of gas remains constant in both the cases

So, "\dfrac{P_1V_1}{RT_1}= \dfrac{P_2V_2}{RT_2}"

So, final volume of the gas will be "V_2=\dfrac{P_1V_1T_2}{T_1P_2}"

Substituting the values "V_2= \dfrac{0.7894\times300\times293}{473\times30}=4.89 L"

Final Volume of the gas is 4.89L